# A Hybrid Quantum-assisted Column Generation Algorithm for the Fleet Conversion Problem

### 11 Jan 2024 v2

@@ 1 Define $R$ such that $R>\max\Gamma_{I}+1$ $\Lambda^{\prime}\leftarrow\emptyset$ while True do
2       Solve RMP and get primal dual couple $(x^{\prime},\mu^{\prime})$ for every model $v$ do
3            Solve SP
@@ 14            Current solution $x^{\prime}$ is the optimal solution
Algorithm 1 Column Generation algorithm for RMP
2.4. Quantum model for the MWIS problem
The MWIS problem can be represented in the form of a QUBO. Our problem in question is the problem SP. For simplicity let $\mu^{\prime}_{k}-\Gamma_{c}^{k}=w_{k}$. The objective function is therefore $\sum\limits_{k}y_{k}w_{k}$. The constraint can be incorporated into the objective function as a penalty (Glover et al., 2019) as follows. The real number $P$ is the penalty strength. Note that the variable $y$ in this section is not related to the other variables $y$ that have appeared previously.
(22) $H=\sum\limits_{k}y_{k}w_{k}-P\bigg{(}\sum\limits_{(i,j)\in E}y_{i}y_{j}\bigg{)}$
@@ 2.4. Quantum model for the MWIS problem
(23) $H=\sum\limits_{k}y_{k}^{2}w_{k}-P\bigg{(}\sum\limits_{(i,j)\in E}y_{i}y_{j}\bigg{)}$
Since $H$ is in the form of a QUBO problem, we can represent it in the following form:
(24) $H=y\mathcal{Q}y^{T}$
where $y$ is the independent set vector and $\mathcal{Q}$ is the QUBO matrix.
As described in Algorithm 3 of (Chatterjee et al., 2023), a QUBO problem can be represented on a quantum computer using only a logarithmic number of qubits. We shall use this algorithm to solve SP. Following are some of the main aspects of the algorithm:
  1. (1)
@@
    (25) $H(\theta)=-\bra{\Psi(\theta)}\mathcal{Q}\ket{\Psi(\theta)}$
    where $\theta=\{\theta_{1}...\theta_{n}\}$ is a set of parameters and |$\ket{\Psi(\theta)}$ is a parameterized ansatz that represents that vector $y$. The interesting thing about this representation is that when $\mathcal{Q}$ is a matrix of size $n\times n$, |$\ket{\Psi}$ is a vector of size $\log n$. Therefore, only $\lceil\log n\rceil$ qubits will be required if we have a problem of size $n$. We put the negative sign as we want to make it a minimization problem.
  2. (2)
    The parameterized state |$\ket{\Psi(\theta)}$ is created as follows:
    (26) |$\ket{\Psi(\theta)}=U(\theta)H^{\otimes n}\ket{0}^{\otimes n}$
    The equation above represents the application of $n$ Hadamard gates on $n$ qubits, all initially in state |$\ket{0}$; followed by the application of a diagonal gate $U(\theta)$ which is of the following form:
    (27) $U(\theta)=\begin{bmatrix}e^{i\pi R(\theta_{1})}&0&0&....\\ 0&e^{i\pi R(\theta_{2})}&0&....\\ ....&....&....&....\\ 0&0&0&e^{i\pi R(\theta_{n})}\end{bmatrix}$
    where
    (28) $R(\theta_{k})=\begin{cases}0\quad&\text{if }\,0\leq\theta_{k}<\pi\\ 1\quad&\text{if }\,\pi\leq\theta_{k}<2\pi\\ \end{cases}$
  3. (3)
    Once we have the ansatz we can measure the expectation value (25) on a quantum computer.
  4. (4)
    Using a classical optimizer such as the genetic algorithm, optimize the parameters $\theta$.
    (29) $H^{*}(\theta^{*})=\min H(\theta)=-\sigma$
@@
    (30) $y_{k}=\frac{1-\exp(i\pi R(\theta_{k}))}{2}=\begin{cases}0\quad&\text{if }\,0\leq\theta_{k}<\pi\\ 1\quad&\text{if }\,\pi\leq\theta_{k}<2\pi\\ \end{cases}$
2.5. Quantum-assisted algorithm to solve the Master Problem
We now have a quantum model to solve the SP. We will call this the Quantum Slave Solver(QSS). To develop a quantum-assisted algorithm, the QSS is used together with a classical slave solver (CSS) to reach the optimal solution. The aim is to use the QSS together with the CSS where the CSS is used only when the QSS is not able to generate a new column.
Figure 2. Progress of the quantum-assisted Algorithm
For every instance, we start the algorithm with the QSS as the slave. If the algorithm is able to generate a new column we use it to solve the master (RMP) and continue with the optimization loop using the QSS. Note that generating a new column here means that the MWIS solution found gives a positive $\sigma$ (as in equation (21)), hence violating (18). In case QSS is not able to generate a new column, we check whether the CSS can generate a new column. If the CSS is able the generate a new column, we solve the RMP using this new column and continue the optimization loop with the QSS. If even the CSS is unable to generate a new column, we have the optimal solution to the problem as the previous solution to the RMP obtained. This is explained diagrammatically in figure 2.
@@ 1 Define $R$ such that $R>\max\Gamma_{I}+1$ $\Lambda^{\prime}\leftarrow\emptyset$ while True do
2       Solve RMP and get primal dual couple $(x^{\prime},\mu^{\prime})$ for every model $v$ do
3            Solve SP using QSS
@@ 26                  Current solution $x^{\prime}$ is the optimal solution
Algorithm 2 Quantum-assisted Column Generation algorithm for RMP
3. Experimental Results
In this section, we demonstrate experiments carried out using tours of sizes 32 and 64. Synthetic data was generated with start and end times of tours as well as sets of allowed vehicles for every tour. There are five different vehicle types and the cardinality of the set of allowed vehicles is 3. For simplicity, we do not generate location data. It can, however be added easily which will only change the density of the incompatibility graph. All runs of the quantum algorithm in this section are done using a quantum simulator.
@@ 3. Experimental Results
Table 1. Results for 32 and 64 Tours.
@@
The number of successful quantum solves is compared to the total number of successful solves. A successful solve means that the QSS or CSS was able to generate a new column. In table 1, the number of quantum solves is shown as a percentage of the total number of solves, averaged over all the instances.
@@
Figure 3 shows in detail a complete run of algorithm 2 for an instance of size 64. The first point of a plot is a gray point. This is a quick greedy solution to the problem and is the starting point of our algorithm. This ensures that we already have a decent solution to start with. As we go forward in the plots, we have blue and red points which signify the RMP solutions obtained using the QSS and the CSS respectively. Finally we have the black point which is the optimal point, where both the QSS and CSS were not able generate a new column.
@@
In the above case, there are 16 iterations where the QSS was successful, 1 where QSS was unsuccessful and CSS was successful and 1 where both QSS and CSS were unsuccessful. Note, however, that for every iteration we called the QSS. Hence, there were a total of 18 QSS calls. Every QSS call requires 5 MWIS solutions (one for each vehicle type) each requiring around 300 expectation value measurements, hence the total number of expectation value measurements required were of the order of $10^{4}$. It was therefore not feasible to run the algorithm on a quantum computer since every expectation value measurement on an IBM quantum computer can take from a few seconds to a few hours, depending upon the wait time (or queue time). In addition we were limited to a size of 64 nodes for simulation since for 128 nodes and above, the simulation time was too long to carry out experiments. Here, the QPU might have significant time advantage over the simulators. However, larger sizes require the calculation of more expectation values in order to get even a feasible solution (an independent set). This further limits experiments on IBM QPUs. Implementation of the algorithm on other QPU systems is a viable road forward for such experiments.
Figure 3. Full run of the algorithm for an instance of 64 tours.
4. Discussion and Conclusion
@@
In this paper we have successfully demonstrated a quantum-assisted algorithm to solve the Fleet Conversion problem. This shows firstly the advantage of having an algorithm that scales logarithmically with the size of the problem. This trait helped us model problems of sizes that are well outside the realm of what other quantum variational algorithms can handle. Notably, our algorithm successfully handles instances that, albeit synthetic, represent real-world, non-trivial industrial problem sizes, which displays the potential utility of leveraging quantum computing for solving complex optimization problems.
@@
Secondly, it demonstrates the possibility of harnessing quantum algorithms in conjunction with classical solvers rather than being in competition with them. If the problem is very resource intensive then instead of the using a full classical optimization, a part of the computation can be transferred to the quantum computer, potentially reducing required computational resources. The availability and efficiency of quantum hardware remain subjects of ongoing research.
@@
Despite this uncertainty, notable advancements in the quantum hardware field in recent years, coupled with the qubit-efficiency of the demonstrated algorithm suggests that quantum-assisted algorithms may hold significant promise in the future.
Acknowledgements.
@@
Y.C. and M.R. acknowledge funding from European Union’s Horizon 2020 research and innovation programme, more specifically the $\bra{\text{NE}}\text{AS}\ket{\text{QC}}$ project under grant agreement No. 951821. A part of the methodology presented in the manuscript is protected by a provisionally patent claim ”Method for optimizing a functioning relative to a set of elements and associated computer program product” submission number EP21306155.9 submitted on 26.8.2021.

### 12 Mar 2024 v3

@@ 1 Define $R$ such that $R>\max\Gamma_{I}+1$ $\Lambda^{\prime}\leftarrow\emptyset$ while True do
2       Solve RCP and get primal dual couple $(x^{\prime},\mu^{\prime})$ for every model $v$ do
3            Solve WP
@@ 14            Current solution $x^{\prime}$ is the optimal solution
Algorithm 1 Column Generation algorithm for RCP
2.4 Quantum model for the MWIS problem
The MWIS problem can be represented in the form of a Quadratic Unconstrained Binary Optimization (QUBO) problem. Our problem in question is the problem WP. In (WP), for simplicity, let $\mu^{\prime}_{k}-\Gamma_{c}^{k}=w_{k}$. The objective function is therefore $\sum\limits_{k}y_{k}w_{k}$. The constraint (WP-1) can be incorporated into the objective function as a penalty [18] as follows. The real number $P$ is the penalty strength. Note that the variable $y$ in this section is not related to the other variables $y$ that have appeared previously.
$\mathcal{C}=\sum\limits_{k}y_{k}w_{k}-P\bigg{(}\sum\limits_{(i,j)\in E}y_{i}y_{j}\bigg{)}$ (22)
@@ 2.4 Quantum model for the MWIS problem
$\mathcal{C}=\sum\limits_{k}y_{k}^{2}w_{k}-P\bigg{(}\sum\limits_{(i,j)\in E}y_{i}y_{j}\bigg{)}$ (23)
Since $\mathcal{C}$ is in the form of a QUBO problem, we can represent it in the following form:
$\mathcal{C}=y^{T}\mathcal{Q}y$ (24)
where $y$ is the vector representation of the independent set and $\mathcal{Q}$ is the QUBO matrix.
As described in Algorithm 3 of [21], a QUBO problem can be represented on a quantum computer using only a logarithmic number of qubits. We shall use this algorithm to solve WP. Following are some of the main aspects of the algorithm. Our main aim here is to represent (24) on a quantum computer.
  1. 1.
    Let the size of the problem be $n$, then the number of qubits required will be $N=\lceil\log_{2}n\rceil$. A $N$-qubit parameterized state |$\ket{\Psi(\theta)}$ is created as follows:
    |$\ket{\Psi(\theta)}=U(\theta)H^{\otimes N}\ket{0}^{\otimes N}$ (25)
    The equation above represents the application of $N$ Hadamard ($H$) gates on $N$ qubits, all initially in state |$\ket{0}$; followed by the application of a diagonal gate $U(\theta)$ which is of the following form:
    $U(\theta)=\begin{bmatrix}e^{i\pi R(\theta_{1})}&0&0&....\\ 0&e^{i\pi R(\theta_{2})}&0&....\\ ....&....&....&....\\ 0&0&0&e^{i\pi R(\theta_{n})}\end{bmatrix}$ (26)
    where
    $R(\theta_{k})=\begin{cases}0\quad&\text{if }\,0\leq\theta_{k}<\pi\\ 1\quad&\text{if }\,\pi\leq\theta_{k}<2\pi\\ \end{cases}$ (27)
    Using (26) and (27) in (25), we have:
    |$\ket{\Psi(\theta)}=\begin{bmatrix}e^{i\pi R(\theta_{1})}\\ e^{i\pi R(\theta_{2})}\\ ......\\ e^{i\pi R(\theta_{n})}\end{bmatrix}$ (28)
    |$\ket{\Psi(\theta)}$ is therefore a vector whose terms belong to the set $\{1,-1\}$.
  2. 2.
    Note that when we defined the QUBO in equation (24), the variables were $y\in\{0,1\}$. Since |$\ket{\Psi(\theta)}\in\{1,-1\}$, we need to define our QUBO matrix using variables $z\in\{1,-1\}$. This new matrix is then called the spin-QUBO or sQUBO as defined in [21]. We shall denote the sQUBO matrix as $\mathcal{Q}^{\prime}$, which is a matrix of size $2n\times 2n$. For details regarding how to generate $\mathcal{Q}^{\prime}$ see Appendix A.
  3. 3.
@@
    $\mathcal{C}(\theta)=-\bra{\Psi(\theta)}\mathcal{Q}^{\prime}\ket{\Psi(\theta)}+K$ (29)
    where $\theta=\{\theta_{1}...\theta_{n}\}$ is a set of parameters and |$\ket{\Psi(\theta)}$ is a parameterized ansatz that represents that vector $y$ and $K$ is a constant. The interesting thing about this representation is that we need only $1+\log n$ qubits to represent $\mathcal{Q}^{\prime}$ of size $2n\times 2n$. Therefore, only $N=1+\lceil\log n\rceil$ qubits will be required if we have a problem of size $n$. We put the negative sign as we want to make it a minimization problem.
  4. 4.
    Measure the expectation value (29) on a quantum computer or a simulator. In order to calculate this expectation value we need to decompose $\mathcal{Q}^{\prime}$ into a sum of Pauli strings (see Appendix B).
  5. 5.
    Using a classical optimizer such as the genetic algorithm (GA), optimize the parameters $\theta$.
    $\mathcal{C}^{*}(\theta^{*})=\min\mathcal{C}(\theta)=-\sigma$ (30)
@@
    $y_{k}=\frac{1-\exp(i\pi R(\theta_{k}^{*}))}{2}=\begin{cases}0\quad&\text{if }\,0\leq\theta_{k}^{*}<\pi\\ 1\quad&\text{if }\,\pi\leq\theta_{k}^{*}<2\pi\\ \end{cases}$ (31)
2.5 Quantum-assisted algorithm to solve the Coordinator Problem
We now have a quantum model to solve the WP. We will call this the Quantum Worker Solver(QWS). To develop a quantum-assisted algorithm, the QWS is used together with a classical worker solver (CWS) to reach the optimal solution. The aim is to use the QWS together with the CWS where the CWS is used only when the QWS is not able to generate a new column.
Figure 2: Progress of the quantum-assisted Algorithm
For every instance, we start the algorithm with the QWS as the worker. If the algorithm is able to generate a new column we use it to solve the coordinator (RCP) and continue with the optimization loop using the QWS. Note that generating a new column here means that the MWIS solution found gives a positive $\sigma$ (as in equation (21)), hence violating (18). In case QWS is not able to generate a new column, we check whether the CWS can generate a new column. If the CWS is able the generate a new column, we solve the RCP using this new column and continue the optimization loop with the QWS. If even the CWS is unable to generate a new column, we have the optimal solution to the problem as the previous solution to the RCP obtained. This is explained diagrammatically in figure 2.
@@ 1 Define $R$ such that $R>\max\Gamma_{I}+1$ $\Lambda^{\prime}\leftarrow\emptyset$ while True do
2       Solve RCP and get primal dual couple $(x^{\prime},\mu^{\prime})$ for every model $v$ do
3            Solve WP using QWS
@@ 26                  Current solution $x^{\prime}$ is the optimal solution
Algorithm 2 Quantum-assisted Column Generation algorithm for RCP
3 Experimental Results
In this section, we demonstrate experiments carried out using tours of sizes 32 and 64. Synthetic data was generated with start and end times of tours as well as sets of allowed vehicles for every tour. We decide to have five different vehicle types and the cardinality of the set of allowed vehicles is 3. For simplicity, we do not generate location data. It can, however be added easily which will only change the density of the incompatibility graph. All runs of the quantum algorithm in this section are done using a quantum simulator.
@@ 3 Experimental Results
Table 1: Results for 32 and 64 Tours.
Instance Size
32 64
Solver used for RCP Gurobi
Solver used for CWS Gurobi
Classical optimizer used for variational loop of QWS Genetic algorithm (python package: geneticalgorithm)
Penalty value 10 20
Average time taken to simulate a complete run of the algorithm on an instance
$\approx 70$ minutes
$\approx 24$ hours
Table 2: Supplementary information regarding the experiments
Genetic Algorithm Parameter Instance Size
32 64
max_num_iteration 50 100
population_size 20 40
mutation_probability 0.1
elit_ratio 0.05
crossover_probability 0.5
parents_portion 0.3
crossover_type uniform
max_iteration_without_improv None
Table 3: Parameters of the Genetic Algorithm used in the experiments
Instance Size No. of Qubits GA Population Size GA Iterations for QWS Estimated Time (hours)
128 8 20 100 108
256 9 20 100 3500
Table 4: Estimated time for simulating instances of 128 and 256 Tours.
@@
In order to evaluate the contribution of the QWS, the number of successful quantum solves is compared to the total number of successful solves. A successful solve means that the QWS or CWS was able to generate a new column. In table 1, the number of quantum solves is shown as a percentage of the total number of solves, averaged over all the instances.
Tables 2 and 3 give some important details regarding the experiments.
@@
Figure 3 shows in detail complete runs of algorithm 2 for instances of size 32 and 64. For every run, the first point of a plot is a gray square. This is a quick greedy solution to the problem and is the starting point of our algorithm. This greedy algorithm colors each node one by one, ensuring that neighboring nodes do not share the same color. This ensures that we already have a decent solution to start with. As we go forward in the plots, we have blue diamonds and red pentagons which signify the RCP solutions obtained using the QWS and the CWS respectively. Finally we have the black hexagon which is the optimal point, where both the QWS and CWS were not able generate a new column. The cost function has been normalized using the formula : (Current Solution)(Optimal LP-feasible Solution)Optimal LP-feasible $\frac{\text{(Current Solution)}-\text{(Optimal LP-feasible Solution)}}{\text{Optimal LP-feasible Solution}}$.
The iterations of the algorithm are different from and not to be confused with the GA iterations of the QWS.
@@
Taking the instance with the blue dash-dotted line for 64 tours as an example, there are 16 iterations where the QWS was successful, 1 where QWS was unsuccessful and CWS was successful and 1 where both QWS and CWS were unsuccessful. Note, however, that for every iteration we called the QWS. Hence, there were a total of 18 QWS calls. Every QWS call requires 5 MWIS solutions (one for each vehicle type) each requiring around 300 expectation value measurements, hence the total number of expectation value measurements required were of the order of $10^{4}$. Even though the size was tractable to represent our problem, it was therefore not feasible to run the algorithm on a quantum computer since every expectation value measurement on an IBM quantum computer can take from a few seconds to a few hours, depending upon the wait time (or queue time). In addition we were limited to a size of 64 nodes for simulation since for 128 nodes and above, the simulation time was too long to carry out experiments. We can see in Table 4 the estimated time that might be required to potentially simulate instances of 128 and 256 tours. These estimates are given considering the population-size and number of GA iterations to be 20 and 100 respectively, and considering that there are 25 algorithm iterations to converge to a solution. Here, the QPU should have significant a time advantage over the simulators.
Figure 3: Full run of the algorithm for instances of 32 and 64 tours.
Finally, figures 4 and 5 show the result of a 32-tour instance. This is specifically the instance denoted by a black solid line in figure 3. Figure 4 shows the incompatibility graph and figure 5 shows the corresponding time-window diagram. The nodes in figure 4 do not refer of any specific location. The tour numbers from 0 to 31 are marked on the time-window bars as well as the nodes. The 5 vehicles used to carry out all the tours are marked using different colors.
Figure 4: Incompatibility graph for a 32-tour instance.
Figure 5: Time-window diagram for a 32-tour instance.
4 Discussion and Conclusion
@@
In this paper we have successfully demonstrated a quantum-assisted algorithm to solve the Fleet Conversion problem. This shows firstly the advantage of having an algorithm that scales logarithmically with the size of the problem. This trait helped us model problems of sizes that are well outside the realm of what other quantum variational algorithms can handle. Notably, our algorithm successfully handles instances that, albeit synthetic, represent realistic, non-trivial industrial problem sizes, which displays the potential utility of leveraging quantum computing for solving complex optimization problems.
@@
Secondly, it demonstrates the possibility of harnessing quantum algorithms in conjunction with classical solvers rather than being in competition with them. If the problem is very resource intensive then instead of the using a full classical optimization, a part of the computation can be transferred to the quantum computer, potentially reducing required computational resources (logarithmic representation).
There remain, however, several drawbacks in the algorithm. Firstly, the algorithm requires the decomposition of the Hamiltonian matrix into a sum of Pauli strings. The cost of doing this using the method used here (as shown in appendix B) increases significantly with an increase in the number of qubits. Reducing the time to decompose the matrix could reduce runtime by a large margin for larger instances.
Secondly, larger instances require a larger number of runs on the genetic algorithm and therefore it quickly becomes infeasible to increase problem size. Even if we decide not to increase the number of genetic algorithm runs, the quality of the QWS solution will degrade resulting in more QWS calls, hence a higher number of iterations (considering we require the same level of performance).
The success of the algorithm depends on the accuracy of the QWS solution. Worse QWS solutions inadvertently lead to more QWS calls. Hence the runtime of the algorithm could be quite sensitive to noise in potential runs on a quantum computer. The indicates that this algorithm would be suitable for small but fault tolerant quantum computers rather than larger ones with higher levels of noise. Despite the fact that we are moving towards a large number of physical qubits in the future, the availability of logical (fault-tolerant) qubits will remain limited. Consequently, this approach will continue to be relevant.
@@
The availability and efficiency of quantum hardware remain subjects of ongoing research. Despite this uncertainty, notable advancements in the quantum hardware field in recent years, coupled with the qubit-efficiency of the demonstrated algorithm suggests that quantum-assisted algorithms may hold significant promise in the future.
acknowledgments
@@
Y.C. and M.R. acknowledge funding from European Union’s Horizon 2020 research and innovation programme, more specifically the $\bra{NE}AS\ket{QC}$ project under grant agreement No. 951821. A part of the methodology used in the manuscript is protected by a provisional patent claim ”Method for optimizing a functioning relative to a set of elements and associated computer program product” submission number EP21306155.9 submitted on 26.8.2021.
conflict of interest
The authors declare no conflict of interest.
[…]
Appendix A Converting QUBO to spin-QUBO
Let $z\in\{1,-1\}$ and $y\in\{0,1\}$, then it is quite easy to verify that $y=\dfrac{1-z}{2}$. Using this in equation (22), we have:
$\mathcal{C}=\sum\limits_{k}w_{k}\dfrac{1-z_{k}}{2}-P\bigg{(}\sum\limits_{(i,j)\in E}\dfrac{1-z_{i}}{2}\dfrac{1-z_{j}}{2}\bigg{)}$ (32)
In the search for optimal values of parameters $z$ we can eliminate the constant terms in $\mathcal{C}$ as they only add a constant shift to the cost function. We can therefore aggregate the constant terms to an overall constant $K$. We can therefore simplify (32) as follows:
$\mathcal{C}$ $=\sum\limits_{k}w_{k}\dfrac{1-z_{k}}{2}-P\bigg{(}\sum\limits_{(i,j)\in E}\dfrac{1-z_{i}}{2}\dfrac{1-z_{j}}{2}\bigg{)}$ (33)
$=\dfrac{1}{2}\sum_{k}w_{k}(1-z_{k})-\dfrac{P}{4}\sum_{\begin{subarray}{c}ij\\ i\neq j\end{subarray}}(1-z_{i}-z_{j}+z_{i}z_{j})$ (36)
$=-\dfrac{1}{2}\sum_{k}w_{k}z_{k}+\dfrac{P}{4}\sum_{\begin{subarray}{c}ij\\ i\neq j\end{subarray}}(z_{i}+z_{j}-z_{i}z_{j})+K$ (39)
The above cannot be represented as a QUBO matrix since it has linear terms which cannot be quadratized since $z_{i}^{2}\neq z_{i}$.
Hence, we need to reformulate the problem. In this reformulation the linear terms are represented in the off-diagonal terms instead of the diagonals. Let us have $2n$ variables $\{z_{1},z_{2}....z_{2n}\}$ where $z_{1}...z_{n}\in\{1,-1\}$ and $z_{n+1}...z_{2n}\in\{1\}$. Then to represent a linear variable $z_{i}$, we can have the term $z_{i}z_{i+n}$ where $z_{i+n}=1$. The equation (39) can be rewritten as follows:
$\mathcal{C}=-\dfrac{1}{2}\sum_{k}w_{k}z_{k}z_{k+n}+\dfrac{P}{4}\sum_{\begin{subarray}{c}ij\\ i\neq j\end{subarray}}(z_{i}z_{i+n}+z_{j}z_{j+n}-z_{i}z_{j})+K$ (40)
This can now be written as:
$\mathcal{C}=z^{T}\mathcal{Q}^{\prime}z+K$ (41)
This is the required $\mathcal{Q}^{\prime}$. It is a matrix of size $2n\times 2n$.
Appendix B Calculating the Expectation value of an observable
Given the observable $\mathcal{O}$, we first need to convert into a sum of tensor products of Pauli strings.
Let $\mathcal{O}$ be a $n\times n$ matrix and $S=\{I,X,Y,Z\}^{N}=\{S_{1},S_{2},S_{3},S_{4}\}^{N}$ be the set of Pauli matrices. We can consider $n$ to be a power of $2$ without any loss of generality. If the size of the Hamiltonian matrix is $n^{\prime}$ which is not a power of $2$, we can easily convert it to a size of $n=2^{\log_{2}(n^{\prime})}$, which is a power of $2$. The extra space in the matrix is filled with 00’s.
This Hamiltonian can now represented on $N=\log_{2}(n)$ qubits. Consider the set $J=\{S_{i_{1}}\otimes S_{i2}...\otimes S_{iN}|i_{1},i_{2}....i_{N}\in\{0,1,2,3\}\}$ which consists of all tensor product combinations of the Pauli matrices.
Then the Hamiltonian can decomposed as:
$\mathcal{O}=\sum\limits_{i=1}^{4^{N}}c_{i}J_{i}$ (42)
where the coefficients are:
$c_{i}=\frac{1}{n}Tr(J_{i}\cdot\mathcal{O})$ (43)
The Hamiltonian therefore becomes:
$\mathcal{O}=\frac{1}{n}\sum\limits_{i=1}^{4^{N}}Tr(J_{i}\cdot\mathcal{O})J_{i}$ (44)
The expectation value becomes a sum of the expectation values of all the terms.
[…]